Pippin Inc. has provided the following data concerning one of the products in its standard cost system. Variable manufacturing overhead is applied to basis of direct labor-hours. Standard Quantity or Hours per Standard Price or Inputs Direct materials Direct labor Variable manufacturing overhead Unit of Output 5.0 grams 0.30 hours 0.30 hours Rate $ 7.00 per gram S 21.30 per hour S 9.60 per hour The company has reported the following actual results for the product for June Actual output Raw materials purchased Actual price of raw materials Raw materials used in production Actual direct labor-hours Actual direct labor rate Actual variable overhead rate 8,500units 48,100grams S 7.70per gram 42,490grams 2,300hours S 21.70per hour S 9.80per hour 38) The raw materials price variance for the month is closest to A) $33,670 U B) S29,743 F C) S33,670 F D) $29,743 U 39) The raw materials quantity variance for the month is closest to A) $70 U B) $77 F C) $77 U D) $70 F 40) The labor rate variance for the month is closest to A) $1,020 U B) S920 U C) $1,0201F D) $920 F
Answer #1
38
Raw materials price variance=48100*(7.7-7)= 33670 U
39
Raw materials quantity variance=7*(42490-8500*5)= 70 F
40
Labor rate variance=2300*(21.7-21.3)= 920 U
41
Labor efficiency variance=21.3*(2300-8500*0.3)= 5325 F
42
Variable overhead rate variance=2300*(9.8-9.6)= 460 U
41
Variable overheadefficiency variance=9.6*(2300-8500*0.3)= 2400 F